OpenStudy (anonymous):

How can i simplify this fraction. The task was to derive N(t) and then use simple algebra to simplify to a form such as N'(t) = #N - #N^2 where # are a mix of variables (a, b, c)

OpenStudy (anonymous):

\[N(t) = \frac{ a }{ b+e^{-ct} }\]

OpenStudy (anonymous):

I then derived it and got \[N'(t) = \frac{ -ace ^{-ct} }{ b^2 + 2be^{-ct} + e^{-2ct} }\]

ganeshie8 (ganeshie8):

"derive" is a wrong word use "differentiate" instead

ganeshie8 (ganeshie8):

recall that the derivative of \(\dfrac{1}{f(t)}\) is \(-\dfrac{1}{f^2(t)}*f'(t)\)

ganeshie8 (ganeshie8):

\[N(t) = \frac{ a }{ b+e^{-ct} } \] differentiate both sides with respect to \(t\) and get \[N'(t) = -\frac{ a }{ (b+e^{-ct})^2 } *(b+e^{-ct})' = -\frac{ a }{ (b+e^{-ct})^2 } * (-ce^{-ct}) \]

ganeshie8 (ganeshie8):

with me so far ?

OpenStudy (phi):

***\[ N'(t) = \frac{ -ace ^{-ct} }{ b^2 + 2be^{-ct} + e^{-2ct} }\] *** that looks ok except I think it should be +ac up top. also, you should keep the bottom in the form (a+b e^(-ct) )^2 because you want to get N out of that mess in other words, write it like this: \[ N' = \frac{a}{b+e^{-ct}} \cdot \frac{a}{b+e^{-ct}} \cdot \frac{c}{a} \cdot e^{-ct} \] the first two expressions are N to get rid of the e^-ct, use the original equation for N, and "solve for e-3t" and plug that into the derivative and simplify.

OpenStudy (anonymous):

@ganeshie8 yep, got that

ganeshie8 (ganeshie8):

see if you can follow the next steps : \[N(t) = \frac{ a }{ b+e^{-ct} } \] \[N'(t) = -\frac{ a }{ (b+e^{-ct})^2 } *(b+e^{-ct})' \\~\\~\\ = -\frac{ a }{ (b+e^{-ct})^2 } * (-ce^{-ct}) \\~\\~\\ = -\frac{ a }{ (b+e^{-ct})^2 } * (-ce^{-ct}) \\~\\ = -\frac{ a^2 }{ (b+e^{-ct})^2 } * \dfrac{(\color{red} {b-b}-ce^{-ct})}{a} \\~\\~\\ = -N^2*(\dfrac{b}{a}-\dfrac{1}{N})\\~\\~\\ =N -\dfrac{b}{a}N^2 \]

OpenStudy (anonymous):

wow. Thank you! I would have never thought to have done that.

OpenStudy (phi):

you should note ganeshie8 made a mistake \[ \frac{b+ce^{-ct}}{a} \] is not 1/N (notice the extra factor of c), rather: \[ \frac{1}{N}=\frac{b+e^{-ct}}{a} \] once you reach \[ N' = \frac{a}{b+e^{-ct}} \cdot \frac{a}{b+e^{-ct}} \cdot \frac{c}{a} \cdot e^{-ct} =N^2 \frac{c}{a} \cdot e^{-ct}\] go back to the original equation \[ N(t) = \frac{ a }{ b+e^{-ct} } \\ b+e^{-ct} = \frac{ a }{N} \\ e^{-ct} = \frac{ a }{N}-b \] and substitute for e^-ct : \[ N' = N^2 \frac{c}{a} \cdot e^{-ct} \\ N' = N^2 \cdot \frac{c}{a} \cdot \left(\frac{ a }{N}-b \right) \] that simplifies to \[ N'= cN - \frac{bc}{a} N^2 \]

OpenStudy (anonymous):

i see your point. i missed the c as well that carried from deriving. thank you! i'm glad you both will get medals in the end. your help was greatly appreciated.